# On a counterexample to Mordell’s Pellian equation conjecture: a non-computer based approach

### v1

@@ 2. Preliminaries
$\left(a_{1}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}^{2}\overline{P}_{3}P_{5}\overline{P}_{6}\overline{P}_{7}P_{43}\overline{P}_{53}$, $\left(a_{2}+\sqrt{d}\right)\mathcal{O}_{K}=P_{3}^{2}\overline{P}_{7}\overline{P}_{9}\overline{P}_{16}\overline{P}_{27}\overline{P}_{46}$,
$\left(a_{3}+\sqrt{d}\right)\mathcal{O}_{K}=P_{3}\overline{P}_{11}\overline{P}_{15}\overline{P}_{39}P_{52}P_{53}$, $\left(a_{4}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{3}P_{7}\overline{P}_{12}P_{37}P_{40}\overline{P}_{41}$,
$\left(a_{5}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{3}P_{3}\overline{P}_{12}P_{39}\overline{P}_{48}P_{51}$, $\left(a_{6}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{3}P_{9}P_{14}\overline{P}_{23}P_{35}P_{47}$,
$\left(a_{7}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}\overline{P}_{3}\overline{P}_{5}\overline{P}_{12}P_{31}\overline{P}_{38}P_{46}$, $\left(a_{8}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{3}^{2}P_{4}P_{5}\overline{P}_{10}\overline{P}_{52}P_{54}$,
$\left(a_{9}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}^{2}P_{6}\overline{P}_{19}P_{39}\overline{P}_{46}P_{48}$, $\left(a_{10}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}P_{14}P_{37}P_{46}\overline{P}_{52}\overline{P}_{55}$,
$\left(a_{11}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{3}P_{6}\overline{P}_{25}\overline{P}_{36}\overline{P}_{37}P_{53}$, $\left(a_{12}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{3}P_{5}P_{17}\overline{P}_{28}\overline{P}_{44}\overline{P}_{48}$,
$\left(a_{13}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}\overline{P}_{3}\overline{P}_{10}P_{15}\overline{P}_{21}P_{25}P_{44}$, $\left(a_{14}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{2}P_{9}P_{23}\overline{P}_{30}\overline{P}_{49}\overline{P}_{53}$,
$\left(a_{15}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}P_{3}\overline{P}_{5}P_{8}\overline{P}_{34}P_{37}P_{55}$, $\left(a_{16}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}^{3}P_{5}P_{18}P_{23}\overline{P}_{33}\overline{P}_{43}$,
$\left(a_{17}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}\overline{P}_{3}^{2}\overline{P}_{30}P_{31}\overline{P}_{34}P_{52}$, $\left(a_{18}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{2}\overline{P}_{21}^{2}P_{33}P_{37}\overline{P}_{45}$,
$\left(a_{19}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}^{2}\overline{P}_{7}P_{8}\overline{P}_{13}\overline{P}_{15}^{2}P_{37}$, $\left(a_{20}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}^{5}\overline{P}_{3}\overline{P}_{4}\overline{P}_{12}\overline{P}_{26}P_{41}$,
$\left(a_{21}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{2}\overline{P}_{18}\overline{P}_{19}P_{29}\overline{P}_{38}P_{52}$, $\left(a_{22}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{4}P_{8}\overline{P}_{20}^{2}P_{28}P_{51}$,
$\left(a_{23}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}P_{5}P_{13}\overline{P}_{18}P_{22}^{2}\overline{P}_{35}$, $\left(a_{24}+\sqrt{d}\right)\mathcal{O}_{K}=P_{3}^{2}\overline{P}_{4}\overline{P}_{8}\overline{P}_{26}\overline{P}_{33}\overline{P}_{50}$,
$\left(a_{25}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{5}P_{9}\overline{P}_{12}P_{18}P_{30}\overline{P}_{54}$, $\left(a_{26}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}P_{3}P_{11}\overline{P}_{12}P_{19}P_{36}\overline{P}_{56}$,
$\left(a_{27}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}\overline{P}_{3}^{2}P_{18}P_{29}P_{33}P_{49}$, $\left(a_{28}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{4}P_{15}\overline{P}_{22}\overline{P}_{23}P_{25}P_{55}$,
$\left(a_{29}+\sqrt{d}\right)\mathcal{O}_{K}=P_{6}P_{7}P_{9}P_{10}P_{13}P_{16}P_{19}$, $\left(a_{30}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}\overline{P}_{4}\overline{P}_{5}\overline{P}_{17}P_{24}P_{30}\overline{P}_{42}$,
$\left(a_{31}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{6}\overline{P}_{16}\overline{P}_{17}P_{21}P_{36}^{2}$, $\left(a_{32}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}\overline{P}_{3}P_{6}P_{13}P_{26}P_{46}P_{52}$,
$\left(a_{33}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{2}P_{3}\overline{P}_{6}\overline{P}_{8}\overline{P}_{19}P_{34}P_{46}$, $\left(a_{34}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}\overline{P}_{7}\overline{P}_{10}^{2}P_{15}\overline{P}_{30}\overline{P}_{36}$,
$\left(a_{35}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{5}P_{11}P_{12}P_{39}P_{46}\overline{P}_{55}$, $\left(a_{36}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}P_{3}\overline{P}_{5}P_{26}P_{28}\overline{P}_{29}\overline{P}_{32}$,
$\left(a_{37}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{4}^{2}\overline{P}_{5}P_{6}P_{12}P_{23}P_{56}$, $\left(a_{38}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{6}P_{10}\overline{P}_{15}^{2}\overline{P}_{39}\overline{P}_{40}$,
$\left(a_{39}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{7}P_{14}\overline{P}_{18}P_{20}\overline{P}_{23}P_{54}$, $\left(a_{40}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}P_{3}P_{5}\overline{P}_{9}P_{25}\overline{P}_{45}\overline{P}_{49}$,
$\left(a_{41}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}^{2}\overline{P}_{6}\overline{P}_{21}\overline{P}_{30}P_{49}\overline{P}_{51}$, $\left(a_{42}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{3}P_{11}P_{13}\overline{P}_{23}P_{39}\overline{P}_{46}$,
$\left(a_{43}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}P_{3}\overline{P}_{7}P_{15}P_{32}P_{38}\overline{P}_{47}$, $\left(a_{44}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{4}\overline{P}_{9}P_{21}P_{23}P_{33}P_{34}$,
$\left(a_{45}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{4}P_{10}\overline{P}_{18}\overline{P}_{28}\overline{P}_{31}P_{32}$, $\left(a_{46}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}^{2}P_{4}^{2}P_{15}P_{21}P_{27}P_{39}$,
$\left(a_{47}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}\overline{P}_{3}\overline{P}_{6}^{2}\overline{P}_{30}P_{40}\overline{P}_{56}$, $\left(a_{48}+\sqrt{d}\right)\mathcal{O}_{K}=P_{4}P_{5}\overline{P}_{28}P_{39}\overline{P}_{49}P_{52}$,
$\left(a_{49}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{3}\overline{P}_{6}\overline{P}_{27}P_{30}P_{46}\overline{P}_{50}$, $\left(a_{50}+\sqrt{d}\right)\mathcal{O}_{K}=P_{3}P_{7}^{2}\overline{P}_{8}P_{9}\overline{P}_{24}P_{47}$,
$\left(a_{51}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{3}\overline{P}_{6}\overline{P}_{16}P_{44}P_{45}\overline{P}_{53}$, $\left(a_{52}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{3}\overline{P}_{5}\overline{P}_{22}\overline{P}_{25}P_{42}P_{46}$,
$\left(a_{53}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}\overline{P}_{3}P_{4}\overline{P}_{5}\overline{P}_{6}\overline{P}_{9}\overline{P}_{13}\overline{P}_{39}$, $\left(a_{54}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{3}\overline{P}_{4}\overline{P}_{7}\overline{P}_{11}P_{12}\overline{P}_{14}P_{47}$,
$\left(a_{55}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{3}\overline{P}_{9}P_{17}P_{28}\overline{P}_{31}P_{39}$, $\left(a_{56}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{3}^{2}\overline{P}_{4}P_{7}\overline{P}_{17}\overline{P}_{47}^{2}$,
$\left(a_{57}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}^{2}\overline{P}_{3}\overline{P}_{22}P_{40}P_{50}\overline{P}_{53}$.
@@
${\rm v}_{P_{1}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{7}+b_{8}+b_{9}+b_{11}+b_{13}+b_{15}+b_{16}+b_{20}+b_{27}+b_{28}+b_{30}+b_{31}+b_{34}+b_{36}+b_{37}+b_{39}+$
$b_{40}+b_{41}+b_{47}+b_{49}+b_{51}+b_{53}+b_{54}+b_{57}=2d_{1}$,
${\rm v}_{P_{2}}\left(z\mathcal{O}_{K}\right)=b_{4}+3b_{5}+b_{7}+2b_{9}+b_{10}+2b_{14}+3b_{16}+2b_{18}+5b_{20}+2b_{21}+b_{23}+b_{26}+b_{27}+2b_{33}+b_{34}+$
$b_{38}+b_{40}+b_{42}+b_{43}+b_{45}+b_{52}=d_{2}$,
${\rm v}_{\overline{P}_{2}}\left(z\mathcal{O}_{K}\right)=2b_{1}+b_{6}+b_{12}+b_{13}+b_{15}+b_{17}+2b_{19}+b_{22}+b_{25}+b_{30}+b_{32}+b_{36}+2b_{41}+b_{44}+2b_{46}+b_{47}+$
$b_{53}+b_{55}+2b_{57}=d_{2}$,

### v2

@@ 2. Preliminaries
$\left(a_{1}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}^{2}\overline{P}_{3}P_{5}\overline{P}_{6}\overline{P}_{7}P_{43}\overline{P}_{53}$, $\left(a_{2}+\sqrt{d}\right)\mathcal{O}_{K}=P_{3}^{2}\overline{P}_{7}\overline{P}_{9}\overline{P}_{16}\overline{P}_{27}\overline{P}_{46}$, $\left(a_{3}+\sqrt{d}\right)\mathcal{O}_{K}=P_{3}\overline{P}_{11}\overline{P}_{15}\overline{P}_{39}P_{52}P_{53}$,
$\left(a_{4}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{3}P_{7}\overline{P}_{12}P_{37}P_{40}\overline{P}_{41}$, $\left(a_{5}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{3}P_{3}\overline{P}_{12}P_{39}\overline{P}_{48}P_{51}$, $\left(a_{6}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{3}P_{9}P_{14}\overline{P}_{23}P_{35}P_{47}$,
$\left(a_{7}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}\overline{P}_{3}\overline{P}_{5}\overline{P}_{12}P_{31}\overline{P}_{38}P_{46}$, $\left(a_{8}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{3}^{2}P_{4}P_{5}\overline{P}_{10}\overline{P}_{52}P_{54}$, $\left(a_{9}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}^{2}P_{6}\overline{P}_{19}P_{39}\overline{P}_{46}P_{48}$,
$\left(a_{10}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}P_{14}P_{37}P_{46}\overline{P}_{52}\overline{P}_{55}$, $\left(a_{11}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{3}P_{6}\overline{P}_{25}\overline{P}_{36}\overline{P}_{37}P_{53}$, $\left(a_{12}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{3}P_{5}P_{17}\overline{P}_{28}\overline{P}_{44}\overline{P}_{48}$,
$\left(a_{13}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}\overline{P}_{3}\overline{P}_{10}P_{15}\overline{P}_{21}P_{25}P_{44}$, $\left(a_{14}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{2}P_{9}P_{23}\overline{P}_{30}\overline{P}_{49}\overline{P}_{53}$, $\left(a_{15}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}P_{3}\overline{P}_{5}P_{8}\overline{P}_{34}P_{37}P_{55}$,
$\left(a_{16}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}^{3}P_{5}P_{18}P_{23}\overline{P}_{33}\overline{P}_{43}$, $\left(a_{17}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}\overline{P}_{3}^{2}\overline{P}_{30}P_{31}\overline{P}_{34}P_{52}$, $\left(a_{18}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{2}\overline{P}_{21}^{2}P_{33}P_{37}\overline{P}_{45}$,
$\left(a_{19}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}^{2}\overline{P}_{7}P_{8}\overline{P}_{13}\overline{P}_{15}^{2}P_{37}$, $\left(a_{20}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}^{5}\overline{P}_{3}\overline{P}_{4}\overline{P}_{12}\overline{P}_{26}P_{41}$, $\left(a_{21}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{2}\overline{P}_{18}\overline{P}_{19}P_{29}\overline{P}_{38}P_{52}$,
$\left(a_{22}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{4}P_{8}\overline{P}_{20}^{2}P_{28}P_{51}$, $\left(a_{23}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}P_{5}P_{13}\overline{P}_{18}P_{22}^{2}\overline{P}_{35}$, $\left(a_{24}+\sqrt{d}\right)\mathcal{O}_{K}=P_{3}^{2}\overline{P}_{4}\overline{P}_{8}\overline{P}_{26}\overline{P}_{33}\overline{P}_{50}$,
$\left(a_{25}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{5}P_{9}\overline{P}_{12}P_{18}P_{30}\overline{P}_{54}$, $\left(a_{26}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}P_{3}P_{11}\overline{P}_{12}P_{19}P_{36}\overline{P}_{56}$, $\left(a_{27}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}\overline{P}_{3}^{2}P_{18}P_{29}P_{33}P_{49}$,
$\left(a_{28}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{4}P_{15}\overline{P}_{22}\overline{P}_{23}P_{25}P_{55}$, $\left(a_{29}+\sqrt{d}\right)\mathcal{O}_{K}=P_{6}P_{7}P_{9}P_{10}P_{13}P_{16}P_{19}$, $\left(a_{30}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}\overline{P}_{4}\overline{P}_{5}\overline{P}_{17}P_{24}P_{30}\overline{P}_{42}$,
$\left(a_{31}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{6}\overline{P}_{16}\overline{P}_{17}P_{21}P_{36}^{2}$, $\left(a_{32}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}\overline{P}_{3}P_{6}P_{13}P_{26}P_{46}P_{52}$, $\left(a_{33}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}^{2}P_{3}\overline{P}_{6}\overline{P}_{8}\overline{P}_{19}P_{34}P_{46}$,
$\left(a_{34}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}\overline{P}_{7}\overline{P}_{10}^{2}P_{15}\overline{P}_{30}\overline{P}_{36}$, $\left(a_{35}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{5}P_{11}P_{12}P_{39}P_{46}\overline{P}_{55}$, $\left(a_{36}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}P_{3}\overline{P}_{5}P_{26}P_{28}\overline{P}_{29}\overline{P}_{32}$,
$\left(a_{37}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{4}^{2}\overline{P}_{5}P_{6}P_{12}P_{23}P_{56}$, $\left(a_{38}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{6}P_{10}\overline{P}_{15}^{2}\overline{P}_{39}\overline{P}_{40}$, $\left(a_{39}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{7}P_{14}\overline{P}_{18}P_{20}\overline{P}_{23}P_{54}$,
$\left(a_{40}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{2}P_{3}P_{5}\overline{P}_{9}P_{25}\overline{P}_{45}\overline{P}_{49}$, $\left(a_{41}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}^{2}\overline{P}_{6}\overline{P}_{21}\overline{P}_{30}P_{49}\overline{P}_{51}$, $\left(a_{42}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{3}P_{11}P_{13}\overline{P}_{23}P_{39}\overline{P}_{46}$,
$\left(a_{43}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}P_{3}\overline{P}_{7}P_{15}P_{32}P_{38}\overline{P}_{47}$, $\left(a_{44}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{4}\overline{P}_{9}P_{21}P_{23}P_{33}P_{34}$, $\left(a_{45}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{4}P_{10}\overline{P}_{18}\overline{P}_{28}\overline{P}_{31}P_{32}$,
$\left(a_{46}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}^{2}P_{4}^{2}P_{15}P_{21}P_{27}P_{39}$, $\left(a_{47}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}\overline{P}_{3}\overline{P}_{6}^{2}\overline{P}_{30}P_{40}\overline{P}_{56}$, $\left(a_{48}+\sqrt{d}\right)\mathcal{O}_{K}=P_{4}P_{5}\overline{P}_{28}P_{39}\overline{P}_{49}P_{52}$,
$\left(a_{49}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{3}\overline{P}_{6}\overline{P}_{27}P_{30}P_{46}\overline{P}_{50}$, $\left(a_{50}+\sqrt{d}\right)\mathcal{O}_{K}=P_{3}P_{7}^{2}\overline{P}_{8}P_{9}\overline{P}_{24}P_{47}$, $\left(a_{51}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{3}\overline{P}_{6}\overline{P}_{16}P_{44}P_{45}\overline{P}_{53}$,
$\left(a_{52}+\sqrt{d}\right)\mathcal{O}_{K}=P_{2}\overline{P}_{3}\overline{P}_{5}\overline{P}_{22}\overline{P}_{25}P_{42}P_{46}$, $\left(a_{53}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}\overline{P}_{3}P_{4}\overline{P}_{5}\overline{P}_{6}\overline{P}_{9}\overline{P}_{13}\overline{P}_{39}$, $\left(a_{54}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}P_{3}\overline{P}_{4}\overline{P}_{7}\overline{P}_{11}P_{12}\overline{P}_{14}P_{47}$,
$\left(a_{55}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{2}P_{3}\overline{P}_{9}P_{17}P_{28}\overline{P}_{31}P_{39}$, $\left(a_{56}+\sqrt{d}\right)\mathcal{O}_{K}=\overline{P}_{3}^{2}\overline{P}_{4}P_{7}\overline{P}_{17}\overline{P}_{47}^{2}$, $\left(a_{57}+\sqrt{d}\right)\mathcal{O}_{K}=P_{1}\overline{P}_{2}^{2}\overline{P}_{3}\overline{P}_{22}P_{40}P_{50}\overline{P}_{53}$.
@@
${\rm v}_{P_{1}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{7}+b_{8}+b_{9}+b_{11}+b_{13}+b_{15}+b_{16}+b_{20}+b_{27}+b_{28}+b_{30}+b_{31}+b_{34}+b_{36}+b_{37}+b_{39}+b_{40}+b_{41}+b_{47}+b_{49}+$
$b_{51}+b_{53}+b_{54}+b_{57}=2d_{1}$,
${\rm v}_{P_{2}}\left(z\mathcal{O}_{K}\right)=b_{4}+3b_{5}+b_{7}+2b_{9}+b_{10}+2b_{14}+3b_{16}+2b_{18}+5b_{20}+2b_{21}+b_{23}+b_{26}+b_{27}+2b_{33}+b_{34}+b_{38}+b_{40}+b_{42}+b_{43}+$
$b_{45}+b_{52}=d_{2}$,
${\rm v}_{\overline{P}_{2}}\left(z\mathcal{O}_{K}\right)=2b_{1}+b_{6}+b_{12}+b_{13}+b_{15}+b_{17}+2b_{19}+b_{22}+b_{25}+b_{30}+b_{32}+b_{36}+2b_{41}+b_{44}+2b_{46}+b_{47}+b_{53}+b_{55}+2b_{57}=d_{2}$,